OpenStudy (anonymous):

verify that is an indentity Sin(90 degree+A)=Cos A

jimthompson5910 (jim_thompson5910):

Sin(90 degree+A)=Cos A sin(90+A) = cos(A) sin(90)cos(A) + cos(90)sin(A) = cos(A) (1)cos(A) + (0)sin(A) = cos(A) cos(A) = cos(A)

OpenStudy (anonymous):

ty